# A perturbative construction of primitive forms from log Landau-Ginzburg mirrors of toric manifolds

### v1

@@ 2.4. Integration and the higher residue pairing
Denote by $\text{Sing}(X)$ the singular part of $X$ and define $D_{i}:=X_{i}\cap\text{Sing}(X)$. Denote by $E_{i}$ the set $\{j:\text{dim}(X_{i}\cap X_{j})=n-1\}$ and write $D_{ij}=X_{i}\cap X_{j}$. Then $D_{i}=\cup_{j\in E_{i}}D_{ij}$ is a union of $(n-1)$-dimensional subspaces of $X_{i}$. Note that $\alpha=(\alpha_{i})\in\mathcal{A}^{p,q}_{\log,c}(X)$ implies that for any $i$, $\alpha_{i}\in\mathcal{A}^{0,q}(X_{i},\Omega_{X_{i}}^{p}(\log D_{i}))$. One can now define the Poincaré residue of $\alpha_{i}$ on $D_{ij}$, and in particular,
@@
¯$\left\langle\bar{\partial}\phi,\psi\right\rangle_{\mathcal{A}}=$ $(-1)^{|\phi|+1}\left\langle\phi,\bar{\partial}\psi\right\rangle_{\mathcal{A}}.$
@@
¯$\left\langle\bar{\partial}\alpha,\beta\right\rangle_{\operatorname{PV}}=$ $-(-1)^{|\alpha|}\left\langle\alpha,\bar{\partial}\beta\right\rangle_{\operatorname{PV}},$
@@
By Theorems 2.16 and 2.18, the trace map and the pairing both descend to cohomologies. Applying Corollary 2.13, we can define on $\mathcal{H}_{f}:=H(\operatorname{PV}_{\log}(X)[[u]],Q_{f})$ a trace map and a pairing: for $\alpha(u),\beta(u)\in\mathcal{H}_{f}$,
@@ 2.5. Grading structure and good basis
where $\frac{\partial}{\partial\tilde{\theta}_{i}}$ is an odd operator satisfying $[\frac{\partial}{\partial\tilde{\theta}_{i}},\tilde{\theta}_{i}]=1$. Also, the operators $z_{i}\frac{\partial}{\partial z_{i}}$, $\bar{z}_{i}\frac{\partial}{\partial\bar{z}_{i}}$ and $d\bar{z}_{i}\wedge\frac{\partial}{\partial\text{d}\bar{z}_{i}}$ act on $\mathcal{C}^{0}(\Delta,\mathcal{A}^{0,*})=\bigoplus_{i=1}^{d}\mathcal{A}^{0,*}(X_{i})$ (on each component $X_{i}$, only non-zero $z_{i}$’s and $\bar{z}_{i}$’s in the summand act) and these actions pass on to the subspaces $\mathcal{A}^{0,*}(X)$ using Proposition A.9, while the operator $\tilde{\theta}_{i}\wedge\frac{\partial}{\partial\tilde{\theta}_{i}}$ acts on $\Theta_{X^{\dagger}/0^{\dagger}}^{*}$.
Since $Q_{f}$ can be written as
$Q_{f}=\sum_{i}^{d}\text{d}\bar{z}_{i}\wedge\frac{\partial}{\partial\bar{z}_{i}}+\sum_{i=1}^{n}\left(g_{i}+u\left(z_{i}\frac{\partial}{\partial z_{i}}+\sum_{l=n+1}^{d}a_{li}z_{l}\frac{\partial}{\partial z_{l}}\right)\right)\frac{\partial}{\partial\tilde{\theta}_{i}},$
@@
The higher residue map $\operatorname{Res}_{f}$ and the higher residue pairing $\operatorname{K}_{f}$ both have weight degree $-2n$.
@@
[…]
(6) if $\text{if }\tau_{i}\text{ is contained in }\sigma_{j},\text{ then }i\leq j.$
@@
Without loss of generality, we can assume that $\gamma_{1}$ is generated by $\rho_{1},\cdots,\rho_{p}$, $\sigma$ by $\rho_{1},\cdots,\rho_{k}$, and $\gamma_{2}$ by $\rho_{1},\cdots,\rho_{q}$ with $1\leq p<k\leq q\leq n$. For $p+1\leq i\leq k$, define $\tilde{\theta}_{i}=\theta_{i}+\sum_{l=n+1}^{d}a_{il}\theta_{l}$, then $\tilde{\theta}_{i}\in\Theta^{1}_{X^{\dagger}/0^{\dagger}}$ by (2). We have
@@
If $\deg(\varphi_{i})+\deg(\varphi_{j})\leq 2n$, then $\operatorname{K}_{f}(\varphi_{i},\varphi_{j})\in\mathbb{C}$ by Lemma 2.19. So we assume that $\deg(\varphi_{i})+\deg(\varphi_{j})>2n$. If $\tau_{i}$ and $\tau_{j}$ are not contained in a common cone of $\Sigma$, then $\operatorname{K}_{f}(\varphi_{i},\varphi_{j})=0$ since on each component of $X$ either $\varphi_{i}$ or $\varphi_{j}$ vanishes. Thus we may further assume that $\tau_{i}$ and $\tau_{j}$ are contained in a common maximal cone $\sigma\in\Sigma$. In this case, the number of common variables of $\varphi_{i}$ and $\varphi_{j}$ is $m\geq 1$. Assuming $z_{k}$ is a common variable, then the Moving Lemma 2.20 tells us that we can replace $z_{k}$ in $\varphi_{i}$ by a linear combination of variables not corresponding to rays in $\sigma$. By neglecting the terms whose variables and that of $\varphi_{j}$ do not correspond to rays in a common cone of $\Sigma$, we reduce to the case that $\varphi^{\prime}_{i}$ and $\varphi^{\prime}_{j}$ has $m-1$ common variables. By inductively reducing the number of common variables, we arrived at the case that $\varphi^{\prime}_{i}$ and $\varphi^{\prime}_{j}$ has no common variables and $\deg(\varphi^{\prime}_{i})+\deg(\varphi^{\prime}_{j})>2n$. In this case, $\operatorname{K}_{f}(\varphi^{\prime}_{i},\varphi^{\prime}_{j})=0$ since the corresponding rays cannot lie in a common cone. ∎
@@ 3.1. Unfolding
$\varphi=\pi\circ\tilde{\varphi}:M\to A_{1}(Y,\mathbb{Z}).$
@@
$P_{\varphi}:=\{(m,\varphi(m)+p):m\in|\Sigma|,p\in P\}\subset M\times P^{gp}.$
@@
where $Q:=\{(m,q\circ\varphi(m)+p):m\in|\Sigma|,p\in\mathbb{N}\}$. Dually, we have
@@
Denote $R:=\mathbb{C}[P],T:=\mathbb{C}[t_{1},\cdots,t_{s}]$ (where $s=\mu-\nu+1$) and $\mathscr{R}:=R\otimes_{\mathbb{C}}T$. We define a family of log LG models parametrized by $\mathscr{S}^{{\dagger}}:=\text{Spec}(\mathscr{R})$ (here the log structure is given by the natural map $P\rightarrow\mathscr{R}$) where the superpotential is given by the function $F=f+\sum_{i:\deg(\varphi_{i})\neq 2}t_{i}\varphi_{i}$. Here $s$ is the number of elements of a basis of $H(\Theta^{\bullet}_{X^{\dagger}/0^{\dagger}},\{f,-\})$ whose weight degree is not $2$.
@@
Then $E_{F}F=2F$ and $[E_{F},Q_{F}]=0$. In particular, $E_{F}$ induces a $\mathbb{Z}$-grading on the hypercohomology $\mathcal{H}_{F,k}:=H^{*}(\Theta_{\mathscr{X}^{\dagger}_{k}/\mathscr{S}^{\dagger}_{k}}^{\bullet}[[u]],\{F,-\}+u\partial)$, which is compatible for different values of $k$.
@@
[…]
@@ 3.2. Gauss-Manin connection and the higher residue pairing
It can be verified that $\nabla_{u\frac{\partial}{\partial u}}$ is well-defined on $\mathcal{H}_{F}$ and $[\nabla_{v},\nabla_{u\frac{\partial}{\partial u}}]=0$. This extended connection is called the extended Gauss-Manin connection. We can equally describe the extended Gauss-Manin connection on $\mathcal{H}_{F,k}$ for each $k$ using the same formulation.
For any $v\in\Theta^{1}_{\hat{\mathscr{S}}^{\dagger}}$, the Gauss-Manin connection $\nabla$ induces a $\mathbb{C}((u))$-linear map $\nabla_{v}^{(0)}$ on $\mathcal{H}_{f,\pm}$, which is called the residue of $\nabla$. Let $\varphi_{1},\cdots,\varphi_{\mu}$ be the good basis in Theorem 2.21.
@@
$(\nabla^{(0)}\varphi)\vdash\frac{\text{d}z_{1}}{z_{1}}\wedge\cdots\wedge\frac{\text{d}z_{n}}{z_{n}}=$ $(\text{d}f+u\partial)\left(z_{1}\cdots z_{k}\frac{\text{d}z_{1}}{z_{1}}\wedge\cdots\wedge\frac{\text{d}z_{n}}{z_{n}}\right)$
@@
$\nabla^{(0)}_{v}\varphi=\sum_{l=n+1}^{d}\frac{\text{d}z^{p_{l}}}{z^{p_{l}}}(v)\cdot z_{l}z_{1}\cdots z_{k},$
@@
Now we define a trace map for sections in $\mathcal{H}_{F}$, which will turn out to be a family version of that for $\mathcal{H}_{f}$. Since $F=f+\sum_{i:\deg(\varphi_{i})\neq 2}t_{i}\varphi_{i}$, we can write
@@
[…]
[$[\{F,-\}_{i},(V_{F})_{i}]=1\quad\text{on }X_{i}-0,$
@@
$\operatorname{Tr}(\alpha):=\sum_{i}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{X_{i}}(T_{\rho}^{u}\alpha\vdash\Omega)\wedge\Omega\in\hat{\mathscr{R}}[[u]],$
@@
$\sum_{i}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{X_{i}}Q_{F}T^{u}_{\rho,i}\varphi\wedge\Omega=0,$
@@
$\sum_{i}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{X_{i}}d(T^{u}_{\rho,i}\varphi\wedge\Omega)=0.$
@@
for any pair of $n$-dimensional components of $X_{i},X_{j}$ such that $D_{ij}=X_{i}\cap X_{j}$ has dimension $n-1$; here the subscripts $i$ and $j$ indicate restriction to the corresponding component of $\hat{\mathscr{X}}$.
@@
As $[Q_{F},E_{F}]=0$, there exists a weight grading on $\mathcal{H}_{F}$. It turns out that $\operatorname{K}_{F}$ is homogeneous with respect to this $\mathbb{Z}$-grading.
@@
The higher residue pairing $\operatorname{K}_{F}$ is homogeneous of weight degree $-2n$.
@@
[…]

### v3

@@ 2.4. Integration and the higher residue pairing
Denote by $\text{Sing}(X)$ the singular part of $X$ and define $D_{i}:=X_{i}\cap\text{Sing}(X)$. Denote by $E_{i}$ the set $\{j:\text{dim}(X_{i}\cap X_{j})=n-1\}$ and write $D_{ij}=X_{i}\cap X_{j}$. Then $D_{i}=\cup_{j\in E_{i}}D_{ij}$ is a union of (n1)(n-1)-dimensional subspaces of $X_{i}$. Note that $\alpha=(\alpha_{i})\in\mathcal{A}^{p,q}_{\log,c}(X)$ implies that for any $i$, $\alpha_{i}\in\mathcal{A}^{0,q}(X_{i},\Omega_{X_{i}}^{p}(\log D_{i}))$. One can now define the Poincaré residue of $\alpha_{i}$ on $D_{ij}$, and in particular,
@@
$\left\langle\bar{\partial}\phi,\psi\right\rangle_{\mathcal{A}}=$ $(-1)^{|\phi|+1}\left\langle\phi,\bar{\partial}\psi\right\rangle_{\mathcal{A}}.$
@@
$\left\langle\bar{\partial}\alpha,\beta\right\rangle_{\operatorname{PV}}=$ $-(-1)^{|\alpha|}\left\langle\alpha,\bar{\partial}\beta\right\rangle_{\operatorname{PV}},$
@@
By Theorems 2.16 and 2.18, the trace map and the pairing both descend to cohomologies. Applying Corollary 2.13, we can define on the (formally completed) Brieskorn lattice
$\mathcal{H}_{f}:=H(\operatorname{PV}_{\log}(X)[[u]],Q_{f})$
a trace map and a pairing: for $\alpha(u),\beta(u)\in\mathcal{H}_{f}$,
@@ 2.5. Grading structure and good basis
where $\frac{\partial}{\partial\tilde{\theta}_{i}}$ is an odd operator satisfying $\frac{\partial}{\partial\tilde{\theta}_{i}},\tilde{\theta}_{i}]=1$. Let $\Delta$ be the defining polytope of the toric manifold $Y$ and $\mathcal{C}^{i}(\Delta,\mathcal{A}^{0,*})$ be the sheaf defined in Appendix A. The operators $z_{i}\frac{\partial}{\partial z_{i}}$, $\bar{z}_{i}\frac{\partial}{\partial\bar{z}_{i}}$ and $d\bar{z}_{i}\wedge\frac{\partial}{\partial\text{d}\bar{z}_{i}}$ act on $\mathcal{C}^{0}(\Delta,\mathcal{A}^{0,*})=\bigoplus_{i=1}^{d}\mathcal{A}^{0,*}(X_{i})$ (on each component $X_{i}$, only non-zero $z_{i}$’s and $\bar{z}_{i}$’s in the summand act) and these actions pass on to the subspaces $\mathcal{A}^{0,*}(X)$ using Proposition A.9, while the operator $\tilde{\theta}_{i}\wedge\frac{\partial}{\partial\tilde{\theta}_{i}}$ acts on $\Theta_{X^{\dagger}/0^{\dagger}}^{*}$.
Using relation (1) to write $\partial=\sum_{i=1}^{d}z_{i}\frac{\partial}{\partial z_{i}}\frac{\text{d}z_{i}}{z_{i}}=\sum_{i=1}^{n}(z_{i}\frac{\partial}{\partial z_{i}}+\sum_{l=n+1}^{d}a_{li}z_{l}\frac{\partial}{\partial z_{l}})\frac{\text{d}z_{i}}{z_{i}}$. Then by contraction with $\Omega$ and equation(6) we have
$Q_{f}=$ $\,\bar{\partial}+\{f,-\}+u\partial$
$=$ $\sum_{i}^{d}\text{d}\bar{z}_{i}\wedge\frac{\partial}{\partial\bar{z}_{i}}+\sum_{i=1}^{n}\left(g_{i}+u\left(z_{i}\frac{\partial}{\partial z_{i}}+\sum_{l=n+1}^{d}a_{li}z_{l}\frac{\partial}{\partial z_{l}}\right)\right)\frac{\partial}{\partial\tilde{\theta}_{i}}.$
@@
The higher residue map $\operatorname{Res}_{f}$ and the higher residue pairing $\operatorname{K}_{f}$ both have weight degree 2n-2n.
@@
Now let us describe a set of distinguished generators of $\mathcal{H}_{f}$. As proved in [18, Section 5.2], we can fix an order $\sigma_{1},\cdots,\sigma_{\mu}$ of $n$-dimensional cones of $\Sigma$ such that if we let $\tau_{i}\subset\sigma_{i},\ 1\leq i\leq\mu$ be the intersection of $\sigma_{i}$ with all cones in $\{\sigma_{j}:j>i,\ \dim(\sigma_{i}\cap\sigma_{j})=n-1\}$, then we have
(7) $\text{if }\tau_{i}\text{ is contained in }\sigma_{j},\text{ then }i\leq j.$
@@
Without loss of generality, we assume that $\gamma_{1},\sigma$ and $\gamma_{2}$ are generated by $\{\rho_{1},\cdots,\rho_{p}\}$, $\{\rho_{1},\cdots,\rho_{k}\}$ and $\{\rho_{1},\cdots,\rho_{q}\}$ respectively, where $1\leq p<k\leq q\leq n$. For $p+1\leq i\leq k$, $\tilde{\theta}_{i}=\theta_{i}+\sum_{l=n+1}^{d}a_{il}\theta_{l}\in\Theta^{1}_{X^{\dagger}/0^{\dagger}}$ by (2), so we have
@@
If $\deg(\varphi_{i})+\deg(\varphi_{j})\leq 2n$, then $\operatorname{K}_{f}(\varphi_{i},\varphi_{j})\in\mathbb{C}$ by Lemma 2.19. So we assume that $\deg(\varphi_{i})+\deg(\varphi_{j})>2n$. If $\tau_{i}$ and $\tau_{j}$ are not contained in a common cone of $\Sigma$, then $\operatorname{K}_{f}(\varphi_{i},\varphi_{j})=0$ since on each component of $X$ either $\varphi_{i}$ or $\varphi_{j}$ vanishes. Thus we may further assume that $\tau_{i}$ and $\tau_{j}$ are contained in a common maximal cone $\sigma\in\Sigma$. In this case, the number of common variables of $\varphi_{i}$ and $\varphi_{j}$ is $m\geq 1$. Assuming $z_{k}$ is a common variable, then the Moving Lemma 2.20 tells us that we can replace $z_{k}$ in $\varphi_{i}$ by a linear combination of variables not corresponding to rays in $\sigma$. By neglecting the terms whose variables and that of $\varphi_{j}$ do not correspond to rays in a common cone of $\Sigma$, we reduce to the case that $\varphi^{\prime}_{i}$ and $\varphi^{\prime}_{j}$ has m1m-1 common variables. By inductively reducing the number of common variables, we arrived at the case that $\varphi^{\prime}_{i}$ and $\varphi^{\prime}_{j}$ has no common variables and $\deg(\varphi^{\prime}_{i})+\deg(\varphi^{\prime}_{j})>2n$. In this case, $\operatorname{K}_{f}(\varphi^{\prime}_{i},\varphi^{\prime}_{j})=0$ since the corresponding rays cannot lie in a common cone. ∎
@@ 3.1. Unfolding
$\varphi=\pi\circ\tilde{\varphi}:N\to A_{1}(Y,\mathbb{Z}).$
@@
$P_{\varphi}:=\{(n,\varphi(n)+p):n\in|\Sigma|,p\in P\}\subset N\times P^{gp}.$
@@
[…]
where $Q:=\{(n,q\circ\varphi(n)+p):n\in|\Sigma|,p\in\mathbb{N}\}$. Therefore we have
${\text{Spec}(\mathbb{C}[Q])\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\text{Spec}(\mathbb{C}[P_{\varphi}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\text{Spec}(\mathbb{C}[\mathbb{N}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\text{Spec}(\mathbb{C}[P]).}$
@@
${(\text{Spec}(\mathbb{C}[Q]),Q)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\text{Spec}(\mathbb{C}[P_{\varphi}]),P_{\varphi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\text{Spec}(\mathbb{C}[\mathbb{N}]),\mathbb{N})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\text{Spec}(\mathbb{C}[P]),P).}$
@@
We arrange the good basis $\varphi_{1},\dots,\varphi_{\mu}$ (in Theorem 2.21) so that $\varphi_{1}=1$ (i.e. of weight degree 0), $\varphi_{2},\dots,\varphi_{\nu}$ are of weight degree $2$, while the remaining $\varphi_{j}$’s are of higher weight degrees. Then $s=\mu-\nu+1$ is the number of basis elements whose weight degree is not $2$. Denote $R:=\mathbb{C}[P],T:=\mathbb{C}[t_{1},\cdots,t_{s}]$ and $\mathscr{R}:=R\otimes_{\mathbb{C}}T$; in the following, $R$ parametrizes the smoothing of the target $X$ while $T$ parametrizes the deformation of the superpotential $f$. We define a family of log LG models parametrized by $\mathscr{S}^{{\dagger}}:=\text{Spec}(\mathscr{R})$ (here the log structure is given by the natural map $P\rightarrow\mathscr{R}$) where the superpotential is given by the function $F=f+t_{1}+\sum_{i=2}^{s}t_{i}\varphi_{\nu+i-1}$.
@@
Then $E_{F}F=2F$ and $E_{F},Q_{F}]=0$. In particular, $E_{F}$ induces a $\mathbb{Z}$-grading on the hypercohomology $\mathcal{H}_{F,k}:=H^{*}(\Theta_{\mathscr{X}^{\dagger}_{k}/\mathscr{S}^{\dagger}_{k}}^{\bullet}[[u]],\{F,-\}+u\partial)$, which is compatible for different values of $k$.
@@
$\mathcal{H}^{F}_{k}:=H(\Omega_{\mathscr{X}^{\dagger}_{k}/\mathscr{S}^{\dagger}_{k}}^{\bullet}[[u]],\text{d}F\wedge+u\partial)$
@@ 3.2. Gauss-Manin connection and the higher residue pairing
It can be verified that $\nabla_{u\frac{\partial}{\partial u}}$ is well-defined on $\mathcal{H}_{F}$ and $\nabla_{v},\nabla_{u\frac{\partial}{\partial u}}]=0$. This extended connection is called the extended Gauss-Manin connection. We can equally describe the extended Gauss-Manin connection on $\mathcal{H}_{F,k}$ for each $k$ using the same formulation.
For any $v\in\Theta^{1}_{\hat{\mathscr{S}}^{\dagger}}$, the Gauss-Manin connection $\nabla$ induces a $\mathbb{C}((u))$-linear map $\nabla_{v}^{(0)}$ on $\mathcal{H}_{f,\pm}$, which is called the residue of $\nabla$. It exists because the Gauss-Manin connection $\nabla$ has log poles and hence nontrivial monodromy groups. Explicitly,
$(\nabla^{(0)}\varphi)\vdash\Omega=\frac{1}{u}(\text{d}f+u\partial)\left(\varphi\vdash\Omega\right).$
Let $\varphi_{1},\cdots,\varphi_{\mu}$ be the good basis in Theorem 2.21.
@@
$(u\nabla^{(0)}\varphi)\vdash\frac{\text{d}z_{1}}{z_{1}}\wedge\cdots\wedge\frac{\text{d}z_{n}}{z_{n}}=$ $(\text{d}f+u\partial)\left(z_{1}\cdots z_{k}\frac{\text{d}z_{1}}{z_{1}}\wedge\cdots\wedge\frac{\text{d}z_{n}}{z_{n}}\right)$
@@
$u\nabla^{(0)}_{v}\varphi=\sum_{l=n+1}^{d}\frac{\text{d}z^{p_{l}}}{z^{p_{l}}}(v)\cdot z_{l}z_{1}\cdots z_{k},$
@@
Now we define a trace map for sections in $\mathcal{H}_{F}$, which will turn out to be a family version of that for $\mathcal{H}_{f}$. Since $F=f+t_{1}+\sum_{i=2}^{s}t_{i}\varphi_{\nu+i-1}$, we can write
@@
For any $i$, let $\{F,-\}_{i}$ be defined using $(\text{d}F)_{i}\wedge$ by conjugating with $\Omega$. Then
$\{F,-\}_{i},(V_{F})_{i}]=1\quad\text{on }X_{i}-0,$
@@
$\operatorname{Tr}(\alpha):=\sum_{i}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{X_{i}}(T_{\rho}^{u}\alpha\vdash\Omega)\wedge\Omega\in\hat{\mathscr{R}}[[u]],$
@@
$\sum_{i}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{X_{i}}Q_{F}T^{u}_{\rho,i}\varphi\wedge\Omega=0,$
@@
$\sum_{i}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{X_{i}}d(T^{u}_{\rho,i}\varphi\wedge\Omega)=0.$
@@
for any pair of $n$-dimensional components of $X_{i},X_{j}$ such that $D_{ij}=X_{i}\cap X_{j}$ has dimension n1n-1; here the subscripts $i$ and $j$ indicate restriction to the corresponding component of $\hat{\mathscr{X}}$.
@@
As $Q_{F},E_{F}]=0$, there exists a weight grading on $\mathcal{H}_{F}$. It turns out that $\operatorname{K}_{F}$ is homogeneous with respect to this $\mathbb{Z}$-grading.
@@
The higher residue pairing $\operatorname{K}_{F}$ is homogeneous of weight degree 2n-2n.
@@
    $(u\frac{\text{d}}{\text{d}u}+n)\operatorname{K}_{F}(s_{1},s_{2})=\operatorname{K}_{F}(\nabla_{u\frac{\text{d}}{\text{d}u}}s_{1},s_{2})+\operatorname{K}_{F}(s_{1},\nabla_{u\frac{\text{d}}{\text{d}u}}s_{2})$;